OpenStudy (anonymous):

The equation (x+6)^2+(y+4)^2=36 models the position and range of the source of a radio signal. Describe the position of the source and the range of the signals

OpenStudy (anonymous):

man they really reach for these word problems don't they? you have a circle

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - h)^2 + (y - k)^2 = r^2 }\) \(r\) is the radius. \((h,k)\) is the center. Lol I'm sorry(and late) in telling this but this is an equation of a circle :P

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

general equation of a circle with radius \((h,k)\) radius \(r\) is ... what parth said

Parth (parthkohli):

Compare the equation you are given with the equation of a circle I gave.

Parth (parthkohli):

Radius is r.... I think you meant center \((h,k)\).

OpenStudy (anonymous):

in your case \(r^2=36\) and so \(r=6\) should be able to find the center

Parth (parthkohli):

Lol it's funny to think how both r and (h,k) are radii :P

OpenStudy (anonymous):

Makes my head spin :O